# Continuous-Time Neural Networks Can Stably Memorize Random Spike Trains

### v1

@@ VI-B The Use
TABLE VI: Effect of the additional regularization by (13) on the eigenvalue condition (20), cf. Table IV. Every row summarizes 100 repetitions of the same experiment.
@@
VII Associative Recall
As expected, memorized spike scores can be recalled by noisy partial prompts. Some pertinent experiments are described below. In these experiments, a fraction $\alpha$ of the neurons does not (or not always) operate in its normal mode: these $\alpha L$ neurons are forced to ignore their input and to produce a noisy version of their (prescribed and) memorized spike train instead. These noisy excitations are created by shifting each spike of the prescribed spike train by random zero-mean Gaussian jitter with variance $\sigma_{\mathrm{s}}^{2}$ while preserving the refractory period $\tau_{0}$ between spikes, as detailed in Appendix C.
TABLE VII: Recall from quiet state with $\alpha=50\%$ forced neurons, which are forced to produce the memorized firings with jitter parameter $\sigma_{\mathrm{s}}/\tau_{0}=0.1$. Each group of three rows summarizes 100 repetitions of the same experiment.
@@ VII Associative Recall
VII-A Recall from Quiet State
In this experiment, the network is prepared as described in Section IV, with parameters as in Table II. The network is started in a state of rest (with no firings and all neuron potentials set to zero), but with a fraction $\alpha$ of neurons forced to provide a noisy version of the memorized content as described above.
After about five periods, the precision (4) and the recall (5) have reached a steady state. In Table VII, the precision and the recall are reported, not only for the whole network but also for the $\alpha L$ forced neurons and the remaining autonomous neurons separately (by adapting the range of the outer sums in (4), (5), and (7) accordingly). For every set of parameters, the experiment is repeated 100 times, each with a new random network and a new random prescribed firing score; the table reports the median, the minimum, and the maximum steady-state precision and recall of all 100 repetitions. Note that the temporal accuracy of the autonomously operating neurons exceeds the temporal accuracy of the excitation.
VII-B Switching Between Memorized Content
In this experiment, the memorized random score of firings consists of two segments $\breve{\boldsymbol{x}}_{R}$ and $\breve{\boldsymbol{x}}_{B}$, each of which is periodic with period $T=50\,\tau_{0}$. A fraction $\alpha$ of neurons (henceforth called the forceable neurons) is sometimes forced to produce noisy memorized contents as in Section VII-A.
@@ VII-B Switching Between Memorized Content
  • (1) Autonomously as in Sec. II.
  • (2) The forceable neurons are forced to produce a noisy version of $\breve{\boldsymbol{x}}_{R}$.
  • (3) The forceable neurons are forced to produce a noisy version of $\breve{\boldsymbol{x}}_{B}$.
@@
VIII Conclusion
We have demonstrated (with numerical experiments) that continuous-time neural networks with random (but fixed) transmission delays can store and autonomously reproduce any given random spike trains up to some maximal length $T_{\text{max}}$, with stable accurate relative timing of all spikes, with probability close to one. Moreover, these experiments suggest that $T_{\text{max}}$ scales at least linearly with the number of synaptic inputs per neuron.
@@ VIII Conclusion
Acknowledgment
The paper builds on the foundations laid by Patrick Murer [18], who also helped to get the present work started. We also wish to thank Hampus Malmberg for guidance and support.
Appendix A The Derivatives (15)
@@
[…]
$=$ $\sum_{n^{\prime}}z_{n,n^{\prime}}(\breve{s}_{n}+\Delta_{n})$ (A.3)
@@
Appendix B The Random Periodic Spike Trains
In this section, we provide the definition of, and the sampling algorithm for, the periodic random spike trains that are used in this paper as described in Section III-B.
As stated in Section III-B, the random spike trains for the different neurons are statistically independent. Therefore, it suffices to consider here only a single spike train (i.e., a single neuron).
@@ Appendix B The Random Periodic Spike Trains
The pertinent math is given in Sections B-A and B-B. The resulting sampling algorithm is given in Section B-C. The expected number of spikes per period is given by (B.13) below.
B-A Poisson Process Basics
@@ B-B The Modification
[…]
$=$ $\tilde{\gamma}\lim_{K\rightarrow\infty}C(K,n)\left(\frac{\lambda T/K}{1-\lambda T/K}\right)^{n}\left(1-\frac{\lambda T}{K}\right)^{K},$ (B.8)
@@
Figure 10: The one-to-one correspondence of configurations in Appendix B-B.
@@ B-C The Sampling Algorithm
    For $n^{\prime}=1,\ldots,n-1$, the firing times are
    $s_{n^{\prime}}=s_{0}+n^{\prime}\tau_{0}+u_{n^{\prime}}$ (B.14)
    (cf. Fig. 10).
@@
Appendix C Adding Spike Jitter Respecting the Refractory Period
@@
Figure 12: Factor graph [19] of the joint probability density function of $S_{0},\ldots,S_{n-1}$ in Appendix C, with $g\triangleq\mathbbm{1}[S_{n^{\prime}}-S_{n^{\prime}-1}>\tau_{0}]$ for $n^{\prime}=1,...,n-1$, and optional additional constraints on $S_{0}$ and $S_{n-1}$ expressed by $g_{0}$ and $g_{n-1}$, respectively.
For the experiments in Section VII, the noisy spike times $s_{0},...,s_{n-1}$ are obtained from the nominal spike times $\breve{s}_{0},...,\breve{s}_{n-1}$ by adding zero-mean Gaussian jitter with variance $\sigma_{s}^{2}$ while maintaining the constraint that $s_{0},...,s_{n-1}$ are separated by at least $\tau_{0}$. A factor graph of the pertinent joint probability density function is shown in Fig. 12.
Sampling from this distribution can be done by Gibbs sampling [20] as follows. [^5] The extension to constrained first and/or last spike, e.g., minimum starting time or periodicity requirement, is straightforward. For $m=1,2,\ldots,M\gg 1$, alternate the following two steps, beginning with $s_{k}^{(0)}=\breve{s}_{k}$ for all $k$:
@@
As always with Gibbs sampling, it is difficult to know how large $M$ needs to be. The experiments in Section VII were done with $M=1000$.

### v2

@@ 6.2 The Use
Table 6: Effect of the additional regularization by (13) on the eigenvalue condition (20), cf. Table 4. Every row summarizes 100 repetitions of the same experiment.
@@
7 Associative Recall
As expected, memorized spike scores can be recalled by noisy partial prompts. Some pertinent experiments are described below. In these experiments, a fraction $\alpha$ of the neurons does not (or not always) operate in its normal mode: these $\alpha L$ neurons are forced to ignore their input and to produce a noisy version of their (prescribed and) memorized spike train instead. These noisy excitations are created by shifting each spike of the prescribed spike train by random zero-mean Gaussian jitter with variance $\sigma_{\mathrm{s}}^{2}$ while preserving the refractory period $\tau_{0}$ between spikes, as detailed in Appendix Appendix C: Adding Spike Jitter Respecting the Refractory Period.
Table 7: Recall from quiet state with $\alpha=50\%$ forced neurons, which are forced to produce the memorized firings with jitter parameter $\sigma_{\mathrm{s}}/\tau_{0}=0.1$. Each group of three rows summarizes 100 repetitions of the same experiment.
@@ 7 Associative Recall
7.1 Recall from Quiet State
In this experiment, the network is prepared as described in Section 4, with parameters as in Table 2. The network is started in a state of rest (with no firings and all neuron potentials set to zero), but with a fraction $\alpha$ of neurons forced to provide a noisy version of the memorized content as described above.
After about five periods, the precision (4) and the recall (5) have reached a steady state. In Table 7, the precision and the recall are reported, not only for the whole network but also for the $\alpha L$ forced neurons and the remaining autonomous neurons separately (by adapting the range of the outer sums in (4), (5), and (7) accordingly). For every set of parameters, the experiment is repeated 100 times, each with a new random network and a new random prescribed firing score; the table reports the median, the minimum, and the maximum steady-state precision and recall of all 100 repetitions. Note that the temporal accuracy of the autonomously operating neurons exceeds the temporal accuracy of the excitation.
7.2 Switching Between Memorized Content
In this experiment, the memorized random score of firings consists of two segments $\breve{\boldsymbol{x}}_{R}$ and $\breve{\boldsymbol{x}}_{B}$, each of which is periodic with period $T=50\,\tau_{0}$. A fraction $\alpha$ of neurons (henceforth called the forceable neurons) is sometimes forced to produce noisy memorized contents as in Section 7.1.
@@ 7.2 Switching Between Memorized Content
  • (1)
    Autonomously as in Sec. 2.
  • (2)
    The forceable neurons are forced to produce a noisy version of $\breve{\boldsymbol{x}}_{R}$.
  • (3)
    The forceable neurons are forced to produce a noisy version of $\breve{\boldsymbol{x}}_{B}$.
@@
Conclusion
We have demonstrated (with numerical experiments) that continuous-time neural networks with random (but fixed) transmission delays can store and autonomously reproduce any given random spike trains up to some maximal length $T_{\mathrm{max}}$, with stable accurate relative timing of all spikes, with probability close to one. Moreover, these experiments suggest that $T_{\mathrm{max}}$ scales at least linearly with the number of synaptic inputs per neuron.
@@ Conclusion
Appendix A: The Derivatives (15)
@@
$z_{\ell(n)}(\breve{s}_{n}+\Delta_{n})$ $=$ $\sum_{n^{\prime}}z_{n,n^{\prime}}(\breve{s}_{n}+\Delta_{n})$ (A.3)
@@
Appendix B: The Random Periodic Spike Trains
In this section, we provide the definition of, and the sampling algorithm for, the periodic random spike trains that are used in this paper as described in Section 3.2.
As stated in Section 3.2, the random spike trains for the different neurons are statistically independent. Therefore, it suffices to consider here only a single spike train (i.e., a single neuron).
@@
The pertinent math is given in Sections B.1 and B.2. The resulting sampling algorithm is given in Section B.3. The expected number of spikes per period is given by (B.13) below.
B.1 Poisson Process Basics
@@ B.2 The Modification
$P(N=n)$ $=$ $\tilde{\gamma}\lim_{K\rightarrow\infty}C(K,n)\left(\frac{\lambda T/K}{1-\lambda T/K}\right)^{n}\left(1-\frac{\lambda T}{K}\right)^{K},$ (B.8)
@@
Figure 10: The one-to-one correspondence of configurations in Appendix B.2.
@@ B.3 The Sampling Algorithm
    For $n^{\prime}=1,\ldots,n-1$, the firing times are $s_{n^{\prime}}=s_{0}+n^{\prime}\tau_{0}+u_{n^{\prime}}$ (cf. Fig. 10).
@@
Appendix C: Adding Spike Jitter Respecting the Refractory Period
@@
Figure 12: Factor graph (Loeliger, 2004) of the joint probability density function of $S_{0},\ldots,S_{n-1}$ with $g\triangleq\mathbbm{1}[S_{n^{\prime}}-S_{n^{\prime}-1}>\tau_{0}]$ for $n^{\prime}=1,...,n-1$, and optional additional constraints on $S_{0}$ and $S_{n-1}$ expressed by $g_{0}$ and $g_{n-1}$, respectively.
For the experiments in Section 7, the noisy spike times $s_{0},...,s_{n-1}$ are obtained from the nominal spike times $\breve{s}_{0},...,\breve{s}_{n-1}$ by adding zero-mean Gaussian jitter with variance $\sigma_{s}^{2}$ while maintaining the constraint that $s_{0},...,s_{n-1}$ are separated by at least $\tau_{0}$. A factor graph of the pertinent joint probability density function is shown in Fig. 12.
Sampling from this distribution can be done by Gibbs sampling (Geman & Geman, 1984) as follows. [^5] The extension to constrained first and/or last spike, e.g., minimum starting time or periodicity requirement, is straightforward. For $m=1,2,\ldots,M\gg 1$, alternate the following two steps, beginning with $s_{k}^{(0)}=\breve{s}_{k}$ for all $k$:
@@
As always with Gibbs sampling, it is difficult to know how large $M$ needs to be. The experiments in Section 7 were done with $M=1000$.
Acknowledgment
The paper builds on the foundations laid by Patrick Murer (Murer, 2022), who also helped to get the present work started. We also wish to thank Hampus Malmberg for guidance and support.